Ysidra paints a room that has 400 square feet of wall space in 2and1/2 hours. At this rate,how long will it take her to paint a room that has 720 square feet of wall space?: Ysidra paints a room that has 400 square feet of wall space in 2and1/2 hours. At this rate,how long will it take her to paint a room that has 720 square feet of wall space?
Answer by checkley77(3654)

400/720=2.5/X
400X=2.5*720
400X=1,800
X=1,800/400
X=4.5 HOURS TO PAINT 720FT^2 ROOM.
PROOF:
400/720=2.5/4.5
1,800=1,800